OpenStudy (anonymous):

If the relative frequency of boys among newborn children in the USA is .513, what’s the probability that exactly 4 of the next 5 births will be boys?

OpenStudy (kropot72):

The relative frequency is an estimate of the probability. Using the binomial distribution the required probability is found from \[P(4\ out\ of\ 5\ births\ are\ boys)=\left(\begin{matrix}5 \\ 4\end{matrix}\right)0.513^{4}\times 0.487\] \[=5\times 0.513^{4}\times 0.487\]

OpenStudy (anonymous):

Thank you

OpenStudy (kropot72):

You're welcome :)